# Resource theories with completely passive states and passive states as free states

### v1

@@ I Introduction
Work extraction processes are of fundamental importance in quantum thermodynamics [1], for instance recently they played a certain role in the context of quantum batteries [2]. For finite thermally isolated quantum systems, the maximum work extractable through a cycle is upper bounded by a free energy that is obtained by minimizing the final energy with the constraint that the von Neumann entropy remains constant [3]. In general, the maximum work is equal to the quantum ergotropy [4] and the states with zero ergotropy are passive. The free energy can be extracted in the limit of an infinite number of copies [5] and the states that are passive for any number of copies are completely passive. In this paper, the maximum work extractable is viewed as a resource, and thus linked to the paradigm of the resource theories (see, e.g., Ref. [6] for the role of resource theories in quantum thermodynamics). For instance, a paradigmatic resource theory view the entanglement as a resource [7]. In this resource theory, states without entaglement are separable states, which are viewed as free states. Furthermore, local operations and classical communication do not generate entanglement, leaving separable states separable, and thus are viewed as free operations. Similarly, other resource theories can be formulated, for instance a resource theory of quantum coherence [8]. Furthermore, a resource theory of quantum thermodynamics has been introduced in Ref. [9], where the free operations are thermal operations and do not increase the nonequilibrium Helmholtz free energy (see, e.g., Ref. [10]). From a microscopic point of view, work can be extracted through thermal operations only when the system interacts with an environment at a certain initial temperature. In contrast, we focus on reversible work extractions, discussing two resource theories where the resources are the free energy and the ergotropy, respectively.
@@ II.1 Resource theories
We can formulate different resource theories depending on what the resource is. For instance, the quantum entanglement can be viewed as a resource, such that the free state are separable states and the free operations are obtained through local operations and classical communication [7]. When the quantum coherence is viewed as a resource, the incoherent states, i.e., the states diagonal with respect a fixed basis, are free states, and the largest set of free operations is formed by the maximally incoherent operations, which are defined such that they map incoherent states into incoherent states [8]. In the resource theory of quantum thermodynamics [10], the free operations are the so-called thermal operations, which are defined as $\Lambda_{th}(\rho)=\text{Tr}_{E}\left\{U\rho\otimes\gamma^{E}_{\beta}U^{\dagger}\right\}$ with unitary operators $U$ that commutate with the total Hamiltonian. In detail, the thermal operations act on the Hilbert space of a system $S$ having Hamiltonian $H$, and are defined by performing a Stinespring dilation with an environment $E$ having Hamiltonian $H_{E}$ and prepared in the Gibbs state $\gamma^{E}_{\beta}=e^{-\beta H_{E}}/Z_{E}$ at the temperature $T=\beta^{-1}$, with $Z_{E}=\text{Tr}\left\{e^{-\beta H_{E}}\right\}$, and with a global unitary operator $U$ such that $[U,H+H_{E}]=0$. The thermal operations preserve the Gibbs state of the system $S$, $\Lambda_{th}(\gamma_{\beta})=\gamma_{\beta}$, thus the Gibbs state $\gamma_{\beta}=e^{-\beta H}/Z$, with $Z=\text{Tr}\left\{e^{-\beta H}\right\}$, is the free state. It results that not all the CPTP maps that preserve the Gibbs state $\gamma_{\beta}$, forming the set of Gibbs-preserving operations, are thermal operations (see, e.g., Ref. [11]). Furthermore, the nonequilibrium Helmholtz free energy defined as $F_{\beta}(\rho)=TS(\rho||\gamma_{\beta})-T\ln Z$ does not increase under the action of the thermal operations, i.e., $F_{\beta}(\Lambda_{th}(\rho))\leq F_{\beta}(\rho)$, so that it is a monotone. The maximum work extractable is given by the difference $\Delta F_{\beta}(\rho)=F_{\beta}(\rho)-F_{\beta}(\gamma_{\beta})=TS(\rho||\gamma_{\beta})$. We note that in this case a work extraction is microscopically performed through an environment $E$, which initially is in the equilibrium state $\gamma^{E}_{\beta}$. In contrast, in the following we focus on a reversible work extraction, such that the system $S$ is thermally isolated. Thus, in order to extract a nonzero average work, the unitary time-evolution operator must not commutate with the Hamiltonian. We aim to define some free operations that do not increase the maximum work extractable, which is viewed as a resource. In particular, we find that in the definition of these free operations some sort of temperature plays a pivotal role.
@@ II.2 Work extraction
from which it follows Eq. (9).
@@ III Completely passive states as free states
where $|\psi_{d}\rangle=\sum_{k}|\epsilon_{k}\rangle/\sqrt{d}$, $\sigma=d/(d-1)(\gamma_{\beta}-\sigma^{\prime}/d)$ and $\sigma^{\prime}$ is any density matrix such that $\sigma\geq 0$. The map in Eq. (18) is such that $\Lambda_{\beta}(\rho)=\gamma_{\beta}$ for any incoherent state $\rho$, i.e., such that $\rho=\Delta(\rho)$. Then, $\Lambda_{\beta}(\gamma_{\beta^{\prime}})=\gamma_{\beta}$ for all $\beta^{\prime}$ and thus $\Lambda_{\beta}\in\mathcal{M}_{cp}$, so that the condition (F,i) is satisfied. Since, the image of the map $\Lambda_{\beta}$ is not a subset of $C$, we get $\eta_{\Lambda}(\gamma)>0$ and Eq. (12) is satisfied only if we choose the parameter $\beta$ of the map such that $\beta^{-1}\leq T(\gamma)/\eta_{\Lambda}(\gamma)$ when $\rho=\gamma\in C$, and thus the condition (F,ii) is not always satisfied.
@@
$D(\rho,\gamma)=T(\gamma)S(\rho||\gamma)\geq 0\,,$ (19)
@@
with $|\alpha_{kl}|\leq 1$ belongs to $\mathcal{O}_{cp}$. In particular, although these maps preserve the temperature of all the free states, they can decrease the free energy. To give an example of a free operation that does not preserve the Gibbs state, we consider the map $\Lambda_{\beta}$ in Eq. (18). We get $\Lambda_{\tilde{\beta}(\rho)}\in\mathcal{O}_{cp}$ where $\tilde{\beta}(\rho)$ is any function such that $\tilde{\beta}(\rho)\geq\beta(\rho)$ when $\rho\in C$. In this case, the map is not linear since the Kraus operators depend on $\rho$ through $\tilde{\beta}(\rho)$. For the limit case of a thermalizing map $\mathcal{T}_{\beta}$ such that $\mathcal{T}_{\beta}(\rho)=\gamma_{\beta}$ for all $\rho$, we get $\eta_{\mathcal{T}_{\beta}}(C)=0$, thus the temperature of the free states can be arbitrarily increased, i.e., $\mathcal{T}_{\beta}\in\mathcal{O}_{cp}$ for any $\beta\geq 0$. We note that the distance $D(\rho,\gamma)$ in Eq. (19) is contractive under the action of all these maps. Similarly, the thermal operations $\Lambda_{th|\beta}$ for a given value of $\beta$, or more generally all the operations that preserve the Gibbs state $\gamma_{\beta}$ for a certain value of $\beta$, belong to $\mathcal{O}_{cp}$ if the value of $\beta$ depends on $\rho$, e.g., it is $\beta=\beta(\rho)$ when $\rho\in C$. Of course, these maps can be not linear since $\beta$ depends on $\rho$. We note that $M_{cp}(\rho)=S(\rho||\gamma_{\hat{\beta}(\rho)})$, where $\hat{\beta}(\rho)$ minimizes the relative entropy $S(\rho||\gamma_{\beta})$ over $\beta\geq 0$, thus $M_{cp}(\rho)$ is related to the nonequilibrium Helmholtz free energy as $M_{cp}(\rho)=\hat{\beta}(\rho)\Delta F_{\hat{\beta}(\rho)}(\rho)$.
@@ IV Passive states as free states
Let us try to extend the results achieved to the case of free states that are passive states. We consider the set $\Pi$ of the passive states, $\Pi=\{P=\sum_{k}p_{k}|\epsilon_{k}\rangle\langle\epsilon_{k}||p_{k}\geq p_{k+1}\}$, as the set of the free states. The set $\mathcal{O}_{p}$ of the free operations is formed by all the CPTP maps $\Lambda$ that satisfy (E,i) $\Lambda(\Pi)\subseteq\Pi$ and (E,ii) $\mathcal{E}(\Lambda(\rho))\leq\mathcal{E}(\rho)$ for all $\rho$. Similarly to the free energy case, the condition (E,ii) implies (E,i). Then, $\mathcal{O}_{p}\subseteq\mathcal{M}_{p}=\{\Lambda|\Lambda(\Pi)\subseteq\Pi\}$ and $M_{p}(\rho)=\min_{P\in\Pi}S(\rho||P)$ is a monotone, since $M_{p}(\Lambda(\rho))\leq M_{p}(\rho)$ for all $\Lambda\in\mathcal{M}_{p}$. We note that the ergotropy is equal to the free energy, $\mathcal{E}(\rho)=F(\rho)$, when $P_{\rho}$ is a completely passive state, i.e., $P_{\rho}=\gamma_{\rho}$. Then, to generalize the results achieved for the free energy to the case of the passive states as free states, we try to define a nonequilibrium temperature for the passive state $P$ such that it is equal to the temperature functional $T(P)$ when $P=\gamma$. Thus, from the monotone $M_{p}(\rho)$, similarly to Eq. (11), we define a functional
@@
where $T(P|\rho)$ is some nonequilibrium temperature for the passive state $P$. Of course there are infinite ways to define the nonequilibrium temperature of a non-Gibbs passive state $P$, such that $T(P)=\beta^{-1}$ when $P=\gamma_{\beta}$. When we perform the minimization in Eq. (29), we can select a certain definition of the nonequilibrium temperature, among all the possible definitions, depending on the state $\rho$, achieving in this way a nonequilibrium temperature $T(P|\rho)$ that depends on $\rho$. The minimum in Eq. (29) can be achieved for a unique passive state $P^{*}$, such that $\Phi(\rho)=T(P^{*}|\rho)S(\rho||P^{*})$. We can choose the nonequilibrium temperature $T(P|\rho)$ given a state $\rho$ such that $P^{*}=P_{\rho}$ and the functional in Eq. (29) is equal to the ergotropy, i.e., $\Phi(\rho)=\mathcal{E}(\rho)$. In detail, we find that $P^{*}=P_{\rho}$ and
@@
Thus, the distance $D(\rho,\sigma)$ is defined such that, for $\sigma=P\in P$ and $P\neq|\epsilon_{1}\rangle\langle\epsilon_{1}|$, it reads
@@
i.e., the distance between $\rho$ and the set $C$ is equal to the sum of the distance between $\rho$ and the set $\Pi$ and the distance between the closest state $P_{\rho}$ and the set $C$. We note that the distance in general is not contractive under the action of $\Lambda\in\mathcal{O}_{cp}$ such that $\eta_{\Lambda}(C)=1$, although it is contractive when we consider only the states $P=\gamma\in C$. This can be proved by considering the dephasing map $\Delta\in\mathcal{O}_{cp}$ such that $\eta_{\Delta}(C)=1$ and the example previously considered to show that $T(\Delta(P)|\Delta(\rho))>T(P|\rho)$. Similarly to Eq. (41), we find that $D(\Delta(\rho),\Delta(P))>D(\rho,P)$ if
$\ln\frac{p_{1}}{p_{3}}\ln\frac{p_{2}}{p_{3}}>\frac{\epsilon_{3}-\epsilon_{1}}{\epsilon_{3}+\epsilon_{2}-2\epsilon_{1}}\ln\frac{p_{1}^{2}}{p_{2}p_{3}}\ln\frac{p_{2}}{4p_{3}}\,,$ (45)
which is always satisfied for $\epsilon_{2}=\epsilon_{1}$ if $P\neq I/d$. In general, for $\Lambda\in\mathcal{O}_{p}$, from Eq. (35) we get $D(\Lambda(\rho),\Lambda(\gamma))\leq D(\rho,\gamma)$ if and only if the bound in Eq. (36) is saturated, i.e.,
@@
with $F(\rho)-F(\Lambda(\rho))\geq 0$ for all $\rho$. Then, $\Lambda\in\mathcal{O}_{p}$ if $F(P_{\rho})-F(P_{\Lambda(\rho)})\leq 0$ for all $\rho$, which is difficult to satisfy, e.g., it is not satisfied if $\Lambda(\rho)\in C$ for any $\rho\notin C$ with not thermal eigenvalues, although $\Lambda\in\mathcal{O}_{p}$. We note that the set $\Pi$ is convex, i.e., $P=\sum_{i}p_{i}P_{i}\in\Pi$ if $P_{i}\in\Pi$ for all $i$, whereas $C$ is not convex for $d>2$. Then, the free operations of the two sets $\mathcal{O}_{p}$ and $\mathcal{O}_{cp}$ can show a very different structure. To illustrate it with an example, we consider that the ergotropy is convex, i.e., $\mathcal{E}(\rho)\leq\sum_{i}p_{i}\mathcal{E}(\rho_{i})$ for any convex combination $\rho=\sum_{i}p_{i}\rho_{i}$. Then, since $\mathcal{E}(\rho)$ is convex, any convex combination $\Lambda=\sum_{i}p_{i}\Lambda_{i}$, with $\Lambda_{i}\in\mathcal{O}_{p}$ for all $i$, belongs to $\mathcal{O}_{p}$, although we can get $\Lambda\notin\mathcal{O}_{cp}$ if $\Lambda_{i}\in\mathcal{O}_{cp}$ for all $i$.

### v3

@@ I Introduction
Work extraction processes are of fundamental importance in quantum thermodynamics [1], for instance recently they played a certain role in the context of quantum batteries [2]. For finite thermally isolated quantum systems, the maximum work extractable through a cycle is upper bounded by a free energy that is obtained by minimizing the final energy with the constraint that the von Neumann entropy remains constant [3]. In general, the maximum work is equal to the quantum ergotropy [4] and the states with zero ergotropy are passive. The free energy can be extracted in the limit of an infinite number of copies [5] and the states that are passive for any number of copies are completely passive. In this paper, the maximum work extractable is viewed as a resource, and thus linked to the paradigm of the resource theories (see, e.g., Ref. [6] for the role of resource theories in quantum thermodynamics). For instance, a paradigmatic resource theory views the entanglement as a resource [7]. In this resource theory, states without entaglement are separable states, which are viewed as free states. Furthermore, local operations and classical communication do not generate entanglement, leaving separable states separable, and thus are viewed as free operations. Similarly, other resource theories can be formulated, for instance a resource theory of quantum coherence [8]. Furthermore, a resource theory of quantum thermodynamics has been introduced in Ref. [9], where the free operations are thermal operations and do not increase the nonequilibrium Helmholtz free energy (see, e.g., Ref. [10]). From a microscopic point of view, work can be extracted through thermal operations only when the system interacts with an environment at a certain initial temperature. In contrast, we focus on reversible work extractions, discussing two resource theories where the resources are the free energy and the ergotropy, respectively.
@@ II.1 Resource theories
We can formulate different resource theories depending on what the resource is. For instance, the quantum entanglement can be viewed as a resource, such that the free state are separable states and the free operations are obtained through local operations and classical communication [7]. When the quantum coherence is viewed as a resource, the incoherent states, i.e., the states diagonal with respect to a fixed basis, are free states, and the largest set of free operations is formed by the maximally incoherent operations, which are defined such that they map incoherent states into incoherent states [8]. In the resource theory of quantum thermodynamics [10], the free operations are the so-called thermal operations, which are defined as $\Lambda_{th}(\rho)=\text{Tr}_{E}\left\{U\rho\otimes\gamma^{E}_{\beta}U^{\dagger}\right\}$ with unitary operators $U$ that commutate with the total Hamiltonian. In detail, the thermal operations act on the Hilbert space of a system $S$ having Hamiltonian $H$, and are defined by performing a Stinespring dilation with an environment $E$ having Hamiltonian $H_{E}$ and prepared in the Gibbs state $\gamma^{E}_{\beta}=e^{-\beta H_{E}}/Z_{E}$ at the temperature $T=\beta^{-1}$, with $Z_{E}=\text{Tr}\left\{e^{-\beta H_{E}}\right\}$, and with a global unitary operator $U$ such that $[U,H+H_{E}]=0$. The thermal operations preserve the Gibbs state of the system $S$, $\Lambda_{th}(\gamma_{\beta})=\gamma_{\beta}$, thus the Gibbs state $\gamma_{\beta}=e^{-\beta H}/Z$, with $Z=\text{Tr}\left\{e^{-\beta H}\right\}$, is the free state. It results that not all the CPTP maps that preserve the Gibbs state $\gamma_{\beta}$, forming the set of Gibbs-preserving operations, are thermal operations (see, e.g., Ref. [11]). Furthermore, the nonequilibrium Helmholtz free energy defined as $F_{\beta}(\rho)=TS(\rho||\gamma_{\beta})-T\ln Z$ does not increase under the action of the thermal operations, i.e., $F_{\beta}(\Lambda_{th}(\rho))\leq F_{\beta}(\rho)$, so that it is a monotone. The maximum work extractable is given by the difference $\Delta F_{\beta}(\rho)=F_{\beta}(\rho)-F_{\beta}(\gamma_{\beta})=TS(\rho||\gamma_{\beta})$. We note that in this case a work extraction is microscopically performed through an environment $E$, which initially is in the equilibrium state $\gamma^{E}_{\beta}$. In contrast, in the following we focus on a reversible work extraction, such that the system $S$ is thermally isolated. Thus, in order to extract a nonzero average work, the unitary time-evolution operator must not commutate with the Hamiltonian. We aim to define some free operations that do not increase the maximum work extractable, which is viewed as a resource. In particular, we find that in the definition of these free operations some sort of temperature plays a pivotal role.
@@ II.2 Work extraction
from which it follows Eq. (9). ∎
@@ III Completely passive states as free states
where $|\psi_{d}\rangle=\sum_{k}|\epsilon_{k}\rangle/\sqrt{d}$, $\sigma=d/(d-1)(\gamma_{\beta}-\sigma^{\prime}/d)$ and $\sigma^{\prime}$ is any density matrix such that $\sigma\geq 0$. The map in Eq. (18) is such that $\Lambda_{\beta}(\rho)=\gamma_{\beta}$ for any incoherent state $\rho$, i.e., for any state diagonal with respect to the energy basis. Then, $\Lambda_{\beta}(\gamma_{\beta^{\prime}})=\gamma_{\beta}$ for all $\beta^{\prime}$ and thus $\Lambda_{\beta}\in\mathcal{M}_{cp}$, so that the condition (F,i) is satisfied. Since, the image of the map $\Lambda_{\beta}$ is not a subset of $C$, we get $\eta_{\Lambda}(\gamma)>0$ and Eq. (12) is satisfied only if we choose the parameter $\beta$ of the map such that $\beta^{-1}\leq T(\gamma)/\eta_{\Lambda}(\gamma)$ when $\rho=\gamma\in C$, and thus the condition (F,ii) is not always satisfied.
@@
$D(\rho,\gamma)=T(\gamma)S(\rho||\gamma)\,,$ (19)
@@
with $|\alpha_{kl}|\leq 1$ belongs to $\mathcal{O}_{cp}$. In particular, although these maps preserve the temperature of all the free states, they can decrease the free energy. To give an example of a free operation that does not preserve the Gibbs state, we consider the map $\Lambda_{\beta}$ in Eq. (18). We get $\Lambda_{\tilde{\beta}(\rho)}\in\mathcal{O}_{cp}$ where $\tilde{\beta}(\rho)$ is any function such that $\tilde{\beta}(\rho)\geq\beta(\rho)$ when $\rho\in C$. In this case, the map is not linear since the Kraus operators depend on $\rho$ through $\tilde{\beta}(\rho)$. We note that the distance $D(\rho,\gamma)$ in Eq. (19) is contractive under the action of all these maps. For the limit case of a thermalizing map $\mathcal{T}_{\beta}$ such that $\mathcal{T}_{\beta}(\rho)=\gamma_{\beta}$ for all $\rho$ (e.g., achieved from the map $\Lambda_{\beta}$ for $\sigma^{\prime}=\gamma_{\beta}$), we get $\eta_{\mathcal{T}_{\beta}}(C)=0$, thus the distance $D(\rho,\gamma)$ is not contractive under the action of the map $\mathcal{T}_{\beta}$. In this case, the temperature of the free states can be arbitrarily increased, i.e., $\mathcal{T}_{\beta}\in\mathcal{O}_{cp}$ for any $\beta\geq 0$. Similarly, the thermal operations $\Lambda_{th|\beta}$ for a given value of $\beta$, or more generally all the operations that preserve the Gibbs state $\gamma_{\beta}$ for a certain value of $\beta$, belong to $\mathcal{O}_{cp}$ if the value of $\beta$ depends on $\rho$, e.g., it is $\beta=\beta(\rho)$ when $\rho\in C$. Of course, these maps can be not linear since $\beta$ depends on $\rho$. We note that $M_{cp}(\rho)=S(\rho||\gamma_{\hat{\beta}(\rho)})$, where $\hat{\beta}(\rho)$ minimizes the relative entropy $S(\rho||\gamma_{\beta})$ over $\beta\geq 0$, thus $M_{cp}(\rho)$ is related to the nonequilibrium Helmholtz free energy as $M_{cp}(\rho)=\hat{\beta}(\rho)\Delta F_{\hat{\beta}(\rho)}(\rho)$.
@@ IV Passive states as free states
Let us try to extend the results achieved to the case of free states that are passive states. We consider the set $\Pi$ of the passive states, $\Pi=\{P=\sum_{k}p_{k}|\epsilon_{k}\rangle\langle\epsilon_{k}||p_{k}\geq p_{k+1}\}$, as the set of the free states. The set $\mathcal{O}_{p}$ of the free operations is formed by all the CPTP maps $\Lambda$ that satisfy (E,i) $\Lambda(\Pi)\subseteq\Pi$ and (E,ii) $\mathcal{E}(\Lambda(\rho))\leq\mathcal{E}(\rho)$ for all $\rho$. Similarly to the free energy case, the condition (E,ii) implies (E,i). Then, $\mathcal{O}_{p}\subseteq\mathcal{M}_{p}=\{\Lambda|\Lambda(\Pi)\subseteq\Pi\}$ and $M_{p}(\rho)=\min_{P\in\Pi}S(\rho||P)$ is a monotone, since $M_{p}(\Lambda(\rho))\leq M_{p}(\rho)$ for all $\Lambda\in\mathcal{M}_{p}$. We note that the ergotropy is equal to the free energy, $\mathcal{E}(\rho)=F(\rho)$, when $P_{\rho}$ is a completely passive state, i.e., $P_{\rho}=\gamma_{\rho}$. Then, to generalize the results achieved for the free energy to the case of the passive states as free states, we try to define a nonequilibrium temperature for the passive state $P$ such that it is equal to the temperature functional $T(P)$ when $P=\gamma\in C$. Thus, from the monotone $M_{p}(\rho)$, similarly to Eq. (11), we define a functional
@@
where $T(P|\rho)$ is some nonequilibrium temperature for the passive state $P$. Of course there are infinite ways to define the nonequilibrium temperature of a non-Gibbs passive state $P$, such that $T(P)=\beta^{-1}$ when $P=\gamma_{\beta}$. When we perform the minimization in Eq. (29), we can select a certain definition of the nonequilibrium temperature, among all the possible definitions, depending on the state $\rho$, achieving in this way a nonequilibrium temperature $T(P|\rho)$ that depends on $\rho$. The minimum in Eq. (29) can be achieved for a unique passive state $P^{*}$, such that $\Phi(\rho)=T(P^{*}|\rho)S(\rho||P^{*})$. Given a state $\rho$, we can choose the nonequilibrium temperature $T(P|\rho)$ such that $P^{*}=P_{\rho}$ and the functional in Eq. (29) is equal to the ergotropy, i.e., $\Phi(\rho)=\mathcal{E}(\rho)$. In detail, we find that $P^{*}=P_{\rho}$ and
@@
Thus, the distance $D(\rho,\sigma)$ is defined such that, for $\sigma=P\in\Pi$ and $P\neq|\epsilon_{1}\rangle\langle\epsilon_{1}|$, it reads
@@
i.e., the distance between $\rho$ and the set $C$ is equal to the sum of the distance between $\rho$ and the set $\Pi$ and the distance between the closest state $P_{\rho}$ and the set $C$. We note that the distance $D(\rho,P)$ in general is not contractive under the action of $\Lambda\in\mathcal{O}_{cp}$ such that $\eta_{\Lambda}(C)=1$, although it is contractive when we consider only the states $P=\gamma\in C$. This can be proved by considering the dephasing map $\Delta\in\mathcal{O}_{cp}$ such that $\eta_{\Delta}(C)=1$ and the example previously considered to show that $T(\Delta(P)|\Delta(\rho))>T(P|\rho)$. Similarly to Eq. (41), we find that $D(\Delta(\rho),\Delta(P))>D(\rho,P)$ if
$\ln\frac{p_{1}}{p_{3}}\ln(p_{2}p_{3})>\frac{\epsilon_{3}-\epsilon_{1}}{\epsilon_{3}+\epsilon_{2}-2\epsilon_{1}}\ln\frac{p_{1}^{2}}{p_{2}p_{3}}\ln(4p_{2}p_{3})\,,$ (45)
which can be satisfied, e.g., for $\epsilon_{2}=\epsilon_{1}$, $p_{1}=0.6$ and $p_{2}=p_{3}=0.2$. In general, for $\Lambda\in\mathcal{O}_{p}$, from Eq. (35) we get $D(\Lambda(\rho),\Lambda(\gamma))\leq D(\rho,\gamma)$ if and only if the bound in Eq. (36) is saturated, i.e.,
@@
with $F(\rho)-F(\Lambda(\rho))\geq 0$ for all $\rho$. Then, $\Lambda\in\mathcal{O}_{p}$ if $F(P_{\rho})-F(P_{\Lambda(\rho)})\leq 0$ for all $\rho$, which is difficult to satisfy, e.g., it is not satisfied if $\Lambda(\rho)\in C$ for any $\rho\notin C$ with not thermal eigenvalues, although $\Lambda\in\mathcal{O}_{p}$. We get that, if $\Lambda\in\mathcal{O}_{cp}$, then $\Lambda\in\mathcal{O}_{p}$ only if
$\eta_{\Lambda}(\gamma)T(\Lambda(\gamma))+\delta_{\Lambda}(\gamma)\leq T(\gamma)\,,\forall\gamma\in C\,,$ (48)
where we defined
$\delta_{\Lambda}(\gamma)=\frac{F(P_{\rho_{\gamma}})-F(P_{\Lambda(\rho_{\gamma})})}{S(\rho_{\gamma}||\gamma)}$ (49)
and $\rho_{\gamma}$ such that
$\max_{\rho}\frac{F(\Lambda(\rho))}{S(\rho||\gamma)}=\frac{F(\Lambda(\rho_{\gamma}))}{S(\rho_{\gamma}||\gamma)}\,.$ (50)
Proof.
To prove it, it is enough to note that Eq. (47) is satisfied if and only if
$F(P_{\rho})-F(P_{\Lambda(\rho)})\leq T(\gamma)S(\rho||\gamma)-F(\Lambda(\rho))$ (51)
for all $\gamma\in C$, which is equivalent to
$\frac{F(P_{\rho})-F(P_{\Lambda(\rho)})}{S(\rho||\gamma)}\leq T(\gamma)-\frac{F(\Lambda(\rho))}{S(\rho||\gamma)}$ (52)
for all $\gamma\in C$. By noting that $\eta_{\Lambda}(\gamma)$ is defined as in Eq. (14), the latter condition is satisfied only if
$\frac{F(P_{\rho_{\gamma}})-F(P_{\Lambda(\rho_{\gamma})})}{S(\rho_{\gamma}||\gamma)}\leq T(\gamma)-\eta_{\Lambda}(\gamma)T(\Lambda(\gamma))$ (53)
for all $\gamma\in C$, which is equivalent to the condition in Eq. (48). ∎
We note that the set $\Pi$ is convex, i.e., $P=\sum_{i}p_{i}P_{i}\in\Pi$ if $P_{i}\in\Pi$ for all $i$, whereas $C$ is not convex for $d>2$. Then, the free operations of the two sets $\mathcal{O}_{p}$ and $\mathcal{O}_{cp}$ can show a very different structure. To illustrate it with an example, we consider that the ergotropy is convex, i.e., $\mathcal{E}(\rho)\leq\sum_{i}p_{i}\mathcal{E}(\rho_{i})$ for any convex combination $\rho=\sum_{i}p_{i}\rho_{i}$. Then, since $\mathcal{E}(\rho)$ is convex, any convex combination $\Lambda=\sum_{i}p_{i}\Lambda_{i}$, with $\Lambda_{i}\in\mathcal{O}_{p}$ for all $i$, belongs to $\mathcal{O}_{p}$, although we can get $\Lambda\notin\mathcal{O}_{cp}$ if $\Lambda_{i}\in\mathcal{O}_{cp}$ for all $i$.